# Metabolic rate beyond the $3/4$ law

### v1

@@ Abstract
This work presents a discrete theoretical model in which basal metabolic rate $B$ is described as a dynamic function of an organism’s ontogenetic stage $n$. Instead of treating $B$ only as a static function of body mass $M$, we adopt the form $B(n)=B_{0}\,M^{\,b(n)}$, in which the effective scaling exponent $b(n)$ varies systematically throughout development. In contrast to classical approaches, such as Kleiber’s empirical law ($B\propto M^{3/4}$) and the continuous fractal model of West–Brown–Enquist (WBE), which assume a constant exponent, the present framework emphasizes how the metabolic scaling relationship itself can evolve over the life cycle of a single individual. The model is inspired by a Fibonacci-based description of growth in discrete stages, leading to analytic expressions for $b(n)$ that connect ontogenetic progression to changes in the scaling between metabolism and mass. In this setting, Kleiber’s constant $B_{0}\approx 70$ kcal/day is reinterpreted as a metabolic anchoring point, linking the classical law $B\approx 70\,M^{3/4}$ to a developmentally explicit formulation. We show that the resulting trajectory $B(n)$ captures, at a conceptual level, how metabolic scaling can shift from strongly sublinear behavior at early stages towards an almost linear regime as $n$ increases, and that the predicted basal rates remain compatible, in order of magnitude, with values reported for mammals of different sizes. In this way, the work offers a unified framework that connects the evolution of $B(n)$ across ontogeny to the recursive organization of biological growth.
@@ 1 Introduction
The metabolic rate is, in essence, the "pace of life" of an animal, the speed at which its body converts food into energy to sustain all vital functions, from the beating of the heart to the movement of the muscles. This pace is not the same for all organisms; it varies in fascinating ways, influenced by factors such as body size, ambient temperature, and lifestyle. While a giant whale burns energy slowly and steadily, a hummingbird in frantic flight has an extremely accelerated metabolic "engine", requiring a constant intake of food. This relationship between body size and energetic expenditure, classically described by Kleiber [1], allows us to uncover how animals adapt to different environments and how they balance the delicate budget between the energy they obtain and the energy they spend simply to exist.
Kleiber’s Law [1] establishes the relationship between basal metabolic rate $B$ and body mass $M$ in animals in the form $B\propto M^{3/4}$, and emerged from a methodological analysis that synthesized experimental data from multiple sources [2]. Max Kleiber compiled precise measurements of oxygen consumption (obtained by indirect calorimetry under basal conditions) in mammals spanning a wide range of body masses, from small rodents to cattle and elephants. When he attempted to relate mass and metabolism directly in a standard linear plot, the data produced a curve that was difficult to interpret, with values for small animals clustered at one end and values for large animals spread out at the other. The key to revealing the underlying pattern was to represent the data on a log–log plot, graphing $\log(B)$ as a function of $\log(M)$. In this representation, the previously scattered experimental points aligned approximately along a straight line. This alignment is the mathematical signature of a power-law (or scaling law) of the form
$B=a\cdot M^{b},$ (1)
where $B$ represents the basal metabolic rate of the animal, $M$ is its total body mass, $a$ is a proportionality constant that indicates the level of metabolic rate for a unit mass (for example, 1 kg), and $b$ is the scaling exponent, the central parameter of the law. Kleiber applied linear regression analysis to determine the slope of the best-fit line. This slope, denoted by $b$ in the logarithmic equation $\log(B)=b\cdot\log(M)+\log(a)$, corresponds exactly to the scaling exponent in the original power-law relation (Eq. 1). The value Kleiber found for this slope was approximately $0.75$, which became known as the three-quarter law. From the linear fit in logarithmic scale, Kleiber was able to determine not only the scaling exponent $b\approx 0.75$, but also the point at which this line crosses the vertical axis. This intercept, when converted back from the logarithmic scale to the arithmetic scale, corresponds to a value of approximately $70$. Thus, the general expression of Kleiber’s Law, derived directly from his empirical analysis for mammals, can be written in a more complete form as
$B\approx 70\,M^{0.75},$
@@
This expression tells us that, on average, each kilogram of a mammal does not "consume" metabolically in strict proportion to its mass, but according to the three-quarter power law. For example, to estimate the basal metabolism of a 20 kg dog, we apply $B\approx 70\times 20^{0.75}\approx 70\times 9.46\approx 662$ kcal/day. Therefore, Kleiber’s analysis provided not only the scaling law (the exponent $b=3/4$), but also the scaling factor ($a\approx 70$), which turned an abstract mathematical relationship into a concrete predictive tool for comparative physiology. It is important to note that the exact value of $a$ may vary slightly depending on the dataset or taxonomic group, but the $3/4$ exponent was, at the time, regarded by Kleiber as approximately universal among mammals.
@@
While Kleiber’s work established that metabolism scales with mass raised to the $3/4$ power through statistical analysis of experimental data, it was the Metabolic Theory of West, Brown and Enquist [3] (WBE), formulated at the end of the same century, that set out to explain why this specific exponent emerges. The WBE model is based on a fundamental physical principle: the need to optimize the distribution of resources (such as nutrients and oxygen) throughout the body of an organism via hierarchical transport networks with fractal properties, such as the circulatory and respiratory systems. The theory argues that, to efficiently supply every cell in a three-dimensional body from a single source (such as the heart or lungs), the architecture of these networks must obey certain principles of scale invariance and space filling. The mathematical solution of this optimization problem, taking into account the fractal geometry of the networks, leads to a theoretical scaling exponent of $b\approx 3/4$. In this way, the WBE model offers a mechanistic and theoretical explanation for the empirical regularity discovered by Kleiber.
@@
Both Kleiber’s empirical law and the theoretical model of West, Brown and Enquist (WBE) represent fundamental advances in describing metabolic scaling in adult organisms. In their classical formulations, both approaches treat the scaling exponent $b\approx 3/4$ as a fixed, asymptotic value, characteristic of a "fully developed" organism. This assumption, however, shows limitations when confronted with empirical data for organisms undergoing growth. During ontogenetic development, from juvenile stages to maturity, systematic deviations from the three-quarter law are observed, since the effective metabolic exponent is not constant and varies in a predictable way over time. Although the conceptual framework of WBE has been extended to describe ontogenetic growth, it still generally assumes a constant metabolic exponent. Thus, both Kleiber’s analysis and the WBE model, by focusing primarily on the relationship between adult individuals of different sizes (an interspecific approach), do not explicitly incorporate the possibility of a dynamic exponent along the life cycle of a single individual. In this respect, these classical models do not fully capture the dynamics of intraspecific scaling during development, leaving a gap precisely in the stages where energy allocation to the construction of new structures (biosynthesis) directly competes with the maintenance of basal functions.
@@
In this context arises the need for models that describe the metabolic exponent $b$ not as a universal constant, but as a variable function $b(n)$ of the organism’s developmental stage $n$. This limitation of classical models is supported by a growing body of literature documenting variation in the metabolic scaling exponent under different conditions. Glazier shows that the supposed universality of the three-quarter law does not hold either among or within species, with exponents that vary consistently as a function of ontogenetic, ecological and physiological factors [4, 5]. This suggests that a more comprehensive theory of metabolic scaling must incorporate the dynamic nature of the exponent $b$ during ontogenetic growth, reconciling changes in the organism’s energetic priorities with the deviations observed from the three-quarter law.
A previous formulation of the Fibonacci-based ontogenetic model, focusing on the estimation of $b(n)$ and on the quantitative comparison with empirical metabolic data, was presented in Cambui (2025) [6]. In the present work, we adopt a complementary perspective: we emphasize the didactic derivation of the expressions for $b(n)$ and for the metabolic rate $B(n)=70\,M^{b(n)}$, as well as the conceptual interpretation of the ontogenetic dynamics predicted by the model.
@@
In light of this evidence, it becomes natural to seek a theoretical framework in which the metabolic exponent $b$ can, in fact, vary throughout development. Building on the discussion above, we propose a discrete model in which metabolic activity emerges from an intrinsic developmental programming logic. In this model, the regulation of metabolism over growth is not described as a continuous and homogeneous process, but as a sequence of successive, discrete stages. At each stage $n$, the organism reorganizes its energetic priorities (for example, between biosynthesis and maintenance), which is reflected in a distinct effective value for the exponent $b(n)$. In this way, ontogenetic progression is understood as a succession of metabolically programmed phases, providing a formal structure capable of explicitly accommodating the variation of the exponent $b$ as a function of the developmental stage $n$.
The model in a nutshell: Organisms do not only differ in size; they also pass through identifiable developmental stages (e.g., infant, juvenile, adult). We propose that the metabolic scaling exponent $b$ is not fixed, but changes with each developmental stage $n$. This change follows a discrete pattern inspired by the Fibonacci sequence, often used to model stepwise growth in nature. The result is a dynamic scaling law, $B(n)=70\,M^{b(n)},$ in which both $M$ and $b(n)$ are interpreted in relation to the organism’s stage of growth. Kleiber’s static $3/4$ law appears as a special case within this more general, stage-dependent framework, for stages (or species) in which $b(n)\approx 3/4$.
@@ 2 From recursive growth to metabolic scaling
Before proceeding, it is important to clarify how the present section relates to our earlier work. A formal derivation of the Fibonacci-based ontogenetic model, including the construction of the recursive growth scheme and the analytical expression for the exponent $b(n)$, was already presented in Cambui (2025) [6], with emphasis on parameter estimation and quantitative comparison with empirical data. Here, we revisit the same mathematical backbone with a different purpose: to reorganize the derivation in a more didactic way and to make explicit how the recursive growth hypothesis leads naturally to a stage-dependent metabolic law of the form $B(n)=B_{0}M^{b(n)}$. The focus is therefore not on proposing a new model, but on clarifying the logical steps that connect recursive mass growth to metabolic scaling.
Based on the recursive property of the Fibonacci sequence, expressed by $F_{n}=F_{n-1}+F_{n-2}$, we propose a biological analogy. If we imagine that the growth of an organism is not continuous, but occurs in discrete, successive stages (such as juvenile and adult phases), in which the mass acquired at each new stage depends on the gains from the two previous stages, we can transpose this mathematical logic to body mass. This leads to a similar recursive relation for the mass at each stage $n$: $M_{n}=M_{n-1}+M_{n-2}$. Solving this recurrence, as is done for the Fibonacci sequence, we obtain that the total mass after $n$ stages grows approximately exponentially, following
$M(n)\sim M_{0}\cdot\phi^{n},$ (2)
@@
At first sight, the expression $M(n)\sim M_{0}\cdot\phi^{n}$ may seem like a bold or even controversial simplification, since it models the growth of the mass of a complex organism using an exponential formula based on the Golden Ratio, a pattern more frequently associated with static forms than with dynamic processes. However, its value lies less in being a literal and precise description of every gram of tissue, and more in functioning as a minimal conceptual model. It captures the central idea that growth occurs in discrete stages $n$, in which each new step amplifies the existing mass by an approximately constant factor related to $\phi$. Solving Eq. (2) for $n$, we arrive at
@@
in which the scaling exponent $b(n)$ is no longer a constant and now depends explicitly on the ontogenetic stage. This formulation establishes a crucial distinction with respect to the classical law $B=B_{0}M^{b}$. In the classical view, the exponent $b$ is a fixed value, such as Kleiber’s $3/4$, which describes an average relationship among adult individuals of different sizes, and mass $M$ plays the role of the only independent variable. In contrast, in the expression $B(n)=B_{0}M^{b(n)}$, the exponent $b(n)$ becomes a variable that depends on the developmental stage. This means that, for a single individual, the rule that relates metabolism and mass changes systematically throughout its growth, from the initial stages of rapid mass gain to maturity. While the classical law answers the question “how does an adult elephant differ from an adult rat?”, the new formulation seeks to answer “how do the efficiency and metabolic priorities of an organism change as it grows?”. In this way, the introduction of the stage $n$ and of the variable exponent $b(n)$ represents a shift from a comparative, static model to a dynamic model that aims to capture the ontogenetic trajectory of metabolism over the life cycle of an individual.
@@
This also captures an important physiological principle: established tissues, even when they are not at maximal functional activity — such as a skeletal muscle at rest, which is not actively contracting, or hepatic parenchyma between detoxification cycles — still require energy to maintain cellular integrity, membrane potential, and routine protein synthesis. This continuous expenditure, essential for maintaining homeostasis and physiological readiness, constitutes the basal maintenance cost. Thus, we have $B(n)\sim F_{n-1}$ (II).
@@ 3 Dynamics of the scaling exponent as a function of stage $n$
To illustrate how the metabolic exponent $b(n)$ varies throughout development, we apply the model formulas directly, in both the simplified and refined versions, to the first ten stages. Table 1 compares the values of $b_{\text{simp}}(n)=(n-1)/n$ and $b_{\text{ref}}(n)$, which includes the correction term $\log\sqrt{5}$. This comparison is particularly useful, as it allows us to check the internal consistency of the model, identify from which point the two formulations begin to agree approximately, and, most importantly, observe the trajectory traced by the exponent, from the more sensitive initial stages to more advanced phases in which a clear pattern emerges.
Table 1: Values of the simplified exponent $b_{\text{simp}}(n)$ and of the refined exponent $b_{\text{ref}}(n)$ as a function of the developmental stage $n$.
$n$
11 22 33 44 55 66 77 88 99 $10$
$b_{\text{simp}}(n)$ $0.000$ $0.500$ $0.667$ $0.750$ $0.800$ $0.833$ $0.857$ $0.875$ $0.889$ $0.900$
$b_{\text{ref}}(n)$ $2.488$ 2.051-2.051 $0.247$ $0.570$ $0.700$ $0.769$ $0.812$ $0.842$ $0.864$ $0.880$
The analysis of the values in Table 1 reveals three important behaviors that characterize and qualify our Fibonacci-based growth model.
First, we observe that for the earliest developmental stages (especially for $n<4$), the refined model, which takes into account the constant term $\log\sqrt{5}$, yields values of $b(n)$ that fluctuate dramatically, even reaching a negative result for $n=2$. This indicates that the more detailed formulation requires interpretative caution in this initial range. Such fluctuations reflect the sensitivity of the model to the initial mass and to the transition between the first terms of the sequence, which have not yet entered the asymptotic regime. In biological practice, this initial phase may be associated with embryonic or larval stages, in which the relationship between mass and metabolism is highly dynamic and influenced by specific factors that go beyond a simple scaling pattern, so that the model should be viewed there only as a qualitative heuristic.
Second, starting from $n=4$, the values of the two versions begin to converge rapidly. The refined exponent stabilizes and begins to closely follow the trajectory of the simplified version. This indicates that, after the first critical stages, the constant term $\log\sqrt{5}$ loses influence and the behavior of the model is dominated by the underlying recursive logic.
@@
Finally, and most significantly, both curves display a clear and consistent trend: as $n$ increases, the exponent $b(n)$ approaches 1 asymptotically. In the theoretical limit $n\to\infty$, this corresponds to a regime in which metabolism tends to scale approximately linearly with mass ($B\propto M^{1}$). For intermediate stages, the values of $b(n)$ pass through the biologically relevant sublinear range (for example, between $\sim 0.7$ and $0.9$), suggesting a possible ontogenetic transition: while young, rapidly growing organisms operate under a stronger efficiency constraint (more pronounced sublinear scaling), more mature organisms approach a regime in which the metabolic cost becomes progressively more proportional to their total mass.

### v2

@@ Abstract
In earlier work, we introduced a discrete Fibonacci-based ontogenetic model in which the metabolic scaling exponent $b(n)$ is treated as a dynamic function of an organism’s developmental stage, and we estimated $b(n)$ for selected mammalian species. In the present article, we revisit this framework with a complementary aim. Rather than proposing new parameter estimates or statistical fits, we provide a didactic, step-by-step reconstruction of the derivation that leads from the recursive growth hypothesis to analytical expressions for the stage-dependent exponent $b(n)$. Building directly on these previously obtained exponents, we then incorporate Kleiber’s classical result into the model by interpreting the constant $70$ in the law $B\approx 70\,M^{3/4}$ (with $B$ denoting basal metabolic rate and $M$ body mass) as a metabolic ”anchoring point”. This yields a stage-dependent basal metabolic rate of the form $B(n)=70\,M^{b(n)}$, which defines an ontogenetic metabolic trajectory linking recursive growth to changes in scaling. We show, at a conceptual level, how this anchored formulation can describe a shift from strongly sublinear behavior at early stages towards an almost linear regime as development proceeds, while still producing basal rates that are compatible, in order of magnitude, with those reported for mammals of different sizes. In this way, the paper offers a self-contained and pedagogical presentation of the model, emphasizing how ontogenetic changes in metabolic rate can be understood through the combined ideas of Fibonacci-like recursion and metabolic anchoring.
@@ 1 Introduction
Generally speaking, we can define metabolic rate as the ”tempo of life”, understood as the speed at which an organism converts resources into energy to sustain all vital functions, from basic cellular processes to the mechanical work of muscles [1]. This tempo is not the same for all organisms, as it varies with factors such as body size, environmental temperature, and lifestyle. While a giant whale burns energy slowly and steadily, a hummingbird in frenzied flight possesses an extremely accelerated metabolic ”engine,” requiring constant food intake [2]. This relationship between body size and energy expenditure, classically described by Kleiber [3], allows us to understand how life adapts to different environments and how the balance between energy acquired and energy needed for existence is maintained.
Kleiber’s law [3] establishes the relationship between basal metabolic rate $B$ and body mass $M$ in the form $B\propto M^{3/4}$, and emerged from a methodological analysis that synthesized experimental data from multiple sources [4]. Max Kleiber compiled precise measurements of oxygen consumption (obtained by indirect calorimetry under basal conditions) in mammals spanning a wide range of body masses, from small rodents to cattle and elephants. When he attempted to relate mass and metabolism directly in a standard linear plot, the data produced a curve that was difficult to interpret, with values for small animals clustered at one end and values for large animals spread out at the other. The key to revealing the underlying pattern was to represent the data on a log–log plot, graphing $\log(B)$ as a function of $\log(M)$. In this representation, the previously scattered experimental points aligned approximately along a straight line. This alignment is the mathematical signature of a power law (or scaling law) of the form
$B=B_{0}\cdot M^{b},$ (1)
@@
While Kleiber’s work established that metabolism scales with mass raised to the $3/4$ power, it was the Metabolic Theory of West, Brown and Enquist [5] (WBE), that set out to explain why this specific exponent emerges. The WBE model is based on the need to optimize the distribution of resources (nutrients and oxygen) throughout the body. This optimization is achieved through hierarchical transport networks with fractal properties, such as the circulatory and respiratory systems. The theory argues that, to efficiently supply every cell in a three-dimensional body from a single source (such as the heart or lungs), the architecture of these networks must obey certain principles of scale invariance and space filling. The mathematical solution of this optimization problem, taking into account the fractal geometry of the networks, leads to a theoretical scaling exponent of $b\approx 3/4$. In this way, the WBE model offers a mechanistic and theoretical explanation for the law of three-quarters established empirically by Kleiber.
@@
It is important to highlight here that both Kleiber’s empirical law and the West, Brown, and Enquist (WBE) theoretical model represent fundamental advances in the description of the metabolic scale in adult organisms. In their formulations, both approaches treat the scaling exponent $b\approx 3/4$ as a fixed, asymptotic value, characteristic of a “fully developed” organism. This assumption, however, shows limitations when confronted with empirical data for organisms undergoing growth. During ontogenetic development, from juvenile stages to maturity, systematic deviations from the three-quarter law are observed, since the effective metabolic exponent is not constant and varies in a predictable way over time. Although the conceptual framework of WBE has been extended to describe ontogenetic growth, it still generally assumes a constant metabolic exponent. Thus, both Kleiber’s analysis and the WBE model, by focusing primarily on the relationship between adult individuals of different sizes (an interspecific approach), do not explicitly incorporate the possibility of a dynamic exponent along the life cycle of a single individual. In this respect, these classical models do not fully capture the dynamics of intraspecific scaling during development, leaving a gap precisely in the stages where energy allocation to the construction of new structures (biosynthesis) directly competes with the maintenance of basal functions.
@@
In this context, there arises the need for models that describe the metabolic exponent $b$ not as a universal constant, but as a variable function $b(n)$ of the organism’s developmental stage $n$. This limitation of classical models is supported by a growing body of literature documenting variation in the metabolic scaling exponent under different conditions. Glazier shows that the supposed universality of the three-quarter law does not hold either among or within species, with exponents that vary consistently as a function of ontogenetic, ecological, and physiological factors [6, 7]. This suggests that a more comprehensive theory of metabolic scaling must incorporate the dynamic nature of the exponent $b$ during ontogenetic growth. In light of this evidence, it becomes natural to seek a theoretical framework in which the metabolic exponent $b$ can, in fact, vary throughout development.
@@
Based on the discussion above, we proposed a discrete model in which metabolic activity emerges from an intrinsic logic of developmental programming [8]. In this model, the regulation of metabolism throughout growth is not described as a continuous and homogeneous process, but as a sequence of successive and discrete stages in which the organism reorganizes its energetic priorities, resulting in a distinct effective value for the exponent $b(n)$. Ontogenetic progression is thus understood as a succession of metabolically programmed phases that establishes a formal framework capable of explicitly accommodating the variation of the exponent $b$ as a function of the developmental stage $n$.
Our proposal [8] for an ontogenetic model was developed upon the foundation of Fibonacci dynamics, establishing a mathematical framework that aligns with the observed patterns of biological growth. In that formulation, the central focus was the analytical derivation of the stage-dependent exponent $b(n)$, its numerical estimation for different mammals, and a systematic comparison with a metabolic database, aiming to test the predictive robustness of the model. In the present article, we adopt a deliberately distinct and complementary perspective. Here, we directly employ the $b(n)$ exponents obtained in that earlier analysis to calculate the corresponding basal metabolic rates via the formula $B(n)=70\,M^{b(n)}$. Our main objective is not to present new fittings or statistical analyses, but to construct, in a didactic and step-by-step manner, the complete derivation that leads from the recursive growth hypothesis to this functional expression. We thus seek to explicate the internal logic of the model, detail the conceptual interpretation of the ontogenetic dynamics it predicts, and consolidate its theoretical framework, offering a self-contained and pedagogical presentation that was beyond the scope of the previous work.
The model in a nutshell: In earlier work, we introduced a discrete ontogenetic model in which organisms do not differ only in size, but also in developmental stage (e.g., infant, juvenile, adult), labeled by an integer $n$. In that framework, the metabolic scaling exponent $b$ is not fixed; instead, it becomes a stage-dependent quantity $b(n)$. The stage dependence $b(n)$ arises from a Fibonacci-inspired, stepwise growth scheme, in which body mass increases recursively from one stage to the next. In the present paper, we take the exponents $b(n)$ obtained in that previous analysis and use them to define a stage-dependent basal metabolic rate of the form $B(n)=70\,M^{b(n)},$ interpreting both $M$ and $b(n)$ in relation to the organism’s stage of growth. Within this anchored formulation, Kleiber’s three-quarter law appears as a special case for stages (or species) in which $b(n)\approx 3/4$, while deviations from this value along development are made explicit by the function $b(n)$.
@@ 2 From recursive growth to metabolic scaling
Before proceeding, it is important to clarify how the present section relates to our earlier work. A formal derivation of the Fibonacci-based ontogenetic model, including the construction of the recursive growth scheme and the analytical expression for the exponent $b(n)$, was already presented in [8], with emphasis on parameter estimation and quantitative comparison with empirical data. Here, we revisit the same mathematical backbone with a different purpose: to reorganize the derivation in a more didactic way and to make explicit how the recursive growth hypothesis leads naturally to a stage-dependent metabolic law of the form $B(n)=B_{0}M^{b(n)}$.
Based on the recursive property of the Fibonacci sequence, expressed by $F_{n}=F_{n-1}+F_{n-2}$, we propose a biological analogy. If we imagine that the growth of an organism is not continuous, but occurs in discrete, successive stages (such as juvenile and adult phases), in which the mass acquired at each new stage depends on the gains from the two previous stages, we can transpose this mathematical logic to body mass. This leads to a similar recursive relation for the mass at each stage $n$: $M_{n}=M_{n-1}+M_{n-2}$. Solving this recurrence, as is done for the Fibonacci sequence, we obtain that the total mass after $n$ stages grows approximately according to a geometric progression, following
$M(n)\sim M_{0}\,\phi^{n},$ (2)
@@
At first sight, the expression $M(n)\sim M_{0}\,\phi^{n}$ may seem like a bold or even controversial simplification, since it models the growth of the mass of a complex organism using a formula based on the golden ratio, a pattern more frequently associated with static forms than with dynamic processes. However, its value lies less in being a literal and precise description of every gram of tissue, and more in functioning as a minimal conceptual model. It captures the central idea that growth occurs in discrete stages $n$, in which each new step amplifies the existing mass by an approximately constant factor related to $\phi$. Solving Eq. (2) for $n$, we arrive at
@@
in which the scaling exponent $b(n)$ is no longer a constant and now depends explicitly on the ontogenetic stage. This formulation establishes a crucial distinction with respect to the classical law $B=B_{0}M^{b}$. In the classical view, the exponent $b$ is a fixed value, such as Kleiber’s $3/4$, which describes an average relationship among adult individuals of different sizes, and mass $M$ plays the role of the only independent variable. In contrast, in the expression $B(n)=B_{0}M^{b(n)}$, the exponent $b(n)$ becomes a variable that depends on the developmental stage. This means that, for a single individual, the rule that relates metabolism and mass changes systematically throughout its growth, from the initial stages of rapid mass gain to maturity. While the classical law answers the question ”how does an adult elephant differ from an adult rat?”, the new formulation seeks to answer ”how do the efficiency and metabolic priorities of an organism change as it grows?”. In this way, the introduction of the stage $n$ and of the variable exponent $b(n)$ represents a shift from a comparative, static model to a dynamic model that aims to capture the ontogenetic trajectory of metabolism over the life cycle of an individual.
@@
This also captures an important physiological principle: established tissues, even when they are not at maximal functional activity, still require energy to maintain cellular integrity, membrane potential, and routine protein synthesis. This continuous expenditure, essential for maintaining homeostasis and physiological readiness, constitutes the basal maintenance cost. Thus, we have $B(n)\sim F_{n-1}$ (II).
@@ 3 Dynamics of the scaling exponent as a function of stage $n$
The analytical expressions for the scaling exponent, Eqs. (9) and (10), transparently determine how $b(n)$ behaves throughout ontogeny. In our previous work [8], we evaluated these formulas numerically and presented continuous plots of $b(n)$ versus $n$, allowing a detailed comparison with empirical data and with classical baselines such as the WBE exponent $b=3/4$. For the purposes of this article, it is sufficient to summarize the qualitative features of these trajectories that are most relevant for interpreting the behavior of the metabolic rate $B(n)$.
First, the refined expression (9), which retains the finite-size correction term $\log\sqrt{5}$, is highly sensitive during the earliest developmental stages. For small values of $n$, the exponent $b(n)$ can deviate significantly from its later trend, even displaying non-monotonic behavior. This is a direct consequence of the discrete, stage-based foundation of the model: at the onset of the Fibonacci-like growth, the constant term $\log\sqrt{5}$ is comparable in magnitude to $n\log\phi$, making the ratio in Eq. (9) strongly influenced by the initial terms of the sequence. Biologically, this mathematical sensitivity suggests that the model should be interpreted with caution when applied to very early ontogenetic phases (e.g., embryonic or larval stages), where the mass–metabolism relationship is known to be governed by highly dynamic and specific developmental processes that a simple scaling law may not fully capture.
Second, as $n$ increases, the situation stabilizes rapidly. Once the organism progresses beyond the most sensitive initial stages, the term $n\log\phi$ dominates both the numerator and denominator of Eq. (9). Consequently, the refined exponent $b(n)$ converges quickly toward the simplified form $b_{\text{simp}}(n)=(n-1)/n$ given by Eq. (10). In practice, from moderate values of $n$ onwards, the two versions yield very similar exponents. This indicates that the detailed logarithmic correction primarily affects the early range of development, while the overall long-term trend is governed by the underlying recursive structure of the model.
@@
Finally, both formulations share a key qualitative feature: the exponent $b(n)$ increases monotonically with $n$ and asymptotically approaches $1$ as $n\to\infty$. This implies that, in the theoretical limit of very advanced developmental stages, metabolism tends to scale almost linearly with body mass, that is, $B\propto M^{1}$. For the intermediate stages relevant to most empirical data, the values of $b(n)$ lie within the well-documented sublinear interval for mammals (approximately between $0.7$ and $0.9$), as discussed in detail in [8]. This pattern naturally suggests an ontogenetic transition in which young, rapidly growing organisms operate under stronger efficiency constraints, characterized by a more pronounced sublinear scaling (lower $b(n)$), while mature organisms progressively approach a regime in which the metabolic cost becomes more directly proportional to their total mass (higher $b(n)$). Thus, the stage dependence of $b(n)$ encapsulates, at the level of the scaling exponent itself, how an organism’s energetic priorities shift throughout its life cycle.